# Minos: Exploiting Cloud Performance Variation with Function-as-a-Service Instance Selection

### v1

@@ Abstract.
Serverless, FaaS, workflows, optimization, resource contention
1. Introduction
Figure 1. FaaS platforms start new function instances on shared worker nodes. While users can not influence which worker node will be used, underutilized nodes offer two benefits: faster execution times and lower cost. Using the Minos system, newly started instances check if they are running on a node with low utilization. If not, they terminate themselves by crashing. This leads to a pool of better-performing instances that are re-used for subsequent invocations, leading to a compound performance increase.
@@ 1. Introduction
Function-as-a-Service (FaaS) offerings have gained increasing popularity in the last years due to their ease of use (Manner, 2023; Bermbach et al., 2017). With FaaS, the cloud platform is responsible for handling most operational tasks, leaving users free to just write the code that should be executed for every incoming event, commonly referred to as functions (Jonas et al., 2019; Manner, 2023). These functions can then be triggered by events in the cloud platform (such as a document upload) or by common transport protocols such as HTTP. In response to incoming events, the cloud platform assigns a function instance (Brooker et al., 2023). If there is no idle already running instance, the platform needs to start a new instance. This process, called a cold start, leads to increased latency for the request as a worker node needs to load the user code and start an isolated environment to run the user code in (Manner et al., 2018). Often, platforms use microVMs (Brooker et al., 2021) or containerization technologies for this (Platfomr, 2020). Users are billed based on the execution time of the invocation, with millisecond accuracy (Manner, 2023). This makes the model cost-efficient for use cases with irregular or infrequent usage patterns, since there is no resource overhead that users need to maintain. As a cloud service hosted on shared infrastructure, functions suffer from performance variations between instances based on the usage patterns of the rest of the platform (Bermbach et al., 2017). For example, as shown in Figure 1, the worker node that the platform selects to start the instance can have different levels of overall utilization. The instance on the worker node with low utilization will benefit from having to share its resources with fewer neighbors. Taking the CPU as example this means that it will have to perform fewer context switches and that it is more likely data can remain cached. These variations also concern most other cloud services. However, FaaS users suffer twice when the variation leads to them having to use slower resources: not only is the execution time longer, but users are also billed more for these slower requests due to the pay-per-use model. In previous work (Schirmer et al., 2023b), we have shown that some FaaS platforms have more than 10% faster performance during the night, when the overall platform is under less load. In this work, we focus on another kind of variation: when starting multiple instances in parallel, some of these instances are going to be faster than other ones. Selectively only running invocations on the fast instances leads to faster and cheaper function executions. To enable this, we present Minos, a system that runs a quick benchmark during every cold start. If the benchmark reveals a particularly slow instance, Minos re-queues the invocation and then terminates the function instance by crashing it. This leads to the platform starting a new instance that might have better performance, or scheduling the invocation on an already running instance. All instances that are already running must have passed the benchmark during their first request (as otherwise, they would have been shut down). This leads to a pool of known-good instances that outperform the average instance, which is especially useful for longer running applications that re-use these instances often. Our evaluation shows that this leads to a state where it is cheaper and faster for users to waste more resources of the platform.
@@
Figure 2. Overview of the process inside a function instance. The first step of the invocation, called prepare, is always executed. If the invocation is a cold start, a benchmark is executed in parallel. Minos then judges whether the result is better than the elysium threshold, in which case nothing happens (and the instance will be re-used in the future). If the instance does not pass, further execution is interrupted, the invocation is re-queued, and the instance is terminated.
2. A System to Exploit Platform Variability
@@
The following sections go into more detail on choosing how much should be terminated (Section 2.1), how to decide on terminating in a distributed manner (Section 2.2), and how the benchmark should work (Section 2.3).
2.1. How much to terminate?
@@
Figure 3. The total cost of a workflow comprises the cost per millisecond of execution ($c_{\textsf{exec}}$) of all terminated ($d_{\textsf{term}}$), successful ($d_{\textsf{pass}}$) benchmarks and the re-used instances ($d_{\textsf{reuse}}$), as well as the cost per invocation ($c_{\textsf{inv}}$) per terminated, passed, and re-used function.
As shown in Figure 3, the total cost of a workflow is influenced by the cost per unit of execution and the cost per invocation. In general, the cost per unit of execution dominates the total cost: For the smallest Google Cloud Functions function with 128 MB Memory, the cost per invocation ($c_{\textsf{inv}}$) is roughly equivalent to 50ms of execution time ($c_{\textsf{exec}}$). For the biggest function with 32 GB Memory, it is less than 3 ms. Other platforms follow a similar pricing pattern. Thus, for all longer-running functions, the additional invocations made by Minos are quickly offset by using faster instances.
As a kind of emergency exit, Minos needs to keep track of how often an invocation has been re-queued already. If an invocation has already caused to too many instance terminations compared to the expected rate, this means the platform performance is unusually slow or Minos is unlucky with the instances it gets from the platform right now. In this case, the function is marked as good without performing the benchmark, preventing infinite loops. For example, if the expected termination rate is $40\%$, there is a ca. $1\%$ chance that an invocation fails five times in a row ($0.4^{5}$) and a less than 1‰ chance that the same invocation experiences 8 terminations in a row.
2.2. Autonomous Termination with the Elysium Threshold
It would be optimal to have a centralized component that receives the benchmarking results from all instances to decide what instances are routed requests. While this would allow for complicated scheduling mechanisms, it is not feasible for the kind of highly scalable and elastic workloads that FaaS platforms enable users to run. Instead, a newly started function instance decides locally whether it is good enough based on just one value: what result in the benchmark is necessary to not terminate itself, i.e., so that the function will continue living and working on the next invocations. We call this value the elysium factor. This is stored as part of the function configuration, so that Minos does not require any outside communication during calls.
@@ 2.2. Autonomous Termination with the Elysium Threshold
2.3. What and how long to benchmark
The kind of benchmark that should be performed and the acceptable elysium threshold heavily depends on the use case. Previous research has shown that FaaS workloads are most often constrained by network or CPU resources (Eismann et al., 2021). In our preliminary experiments, CPU performance showed the highest variability and was the bottleneck in most applications.
The benchmark itself should ideally not interfere with the normal execution of the function, i.e., it should run while the resource being benchmarked is otherwise unused. The general approach is shown in Figure 2. During a first step of the execution, when resources need to be set up and downloaded, the benchmark and subsequent judging is run in parallel. For example, in a data processing workflow, the raw data needs to be downloaded and loaded into memory. At the same time, Minos can run its benchmark.
@@ 2.3. What and how long to benchmark
3. Evaluation
To demonstrate the feasibility of Minos and evaluate potential performance gains we implement a proof of concept prototype and use it to run a data processing workload.
3.1. Environment Configuration
@@
The use case in this evaluation is using linear regression analysis to predict weather. To achieve this, the function first downloads a CSV file containing weather data for a specific location from previous days. Afterwards, the data is analyzed to predict tomorrow’s weather. While downloading the file, Minos can run its CPU benchmark as the download is network-limited. The following linear regression analysis is CPU intensive, benefitting from the instance selection performed in the previous step. To measure the capability of the CPU, we use matrix multiplication as benchmark in Minos (Werner et al., 2018).
@@
One experiment comprises ten virtual users (VU) sending a request and then waiting one second before sending the next request for 30 minutes. We repeat this experiment every day at the same time for one week. [^2] From 2025-02-03 to 2025-02-09, between 3 and 4 pm UTC As a baseline, we run the same workload at the same time using another function that is exactly the same, except that Minos is completely disabled. All experiments were executed in the europe-west-3 region of Google Cloud Platform with functions configured to have 256 MB RAM (leading to 0.167 vCPU).
@@
For both the baseline and experiment conditions, we measure the time of the function execution, the duration of the benchmark execution, whether it was successful, how often the invocation has been re-tried (retry count), how long the downloading step took (download duration), and how long the linear regression took (analysis duration).
3.2. Results
@@
Figure 4. The linear regression step was, on average, faster every day the experiment was repeated. The maximum improvement was $>13\%$ on day two, and minimum improvement $4.3\%$ on day three and five. Note that the y-axis is limited to values between 1,000 and 3,000.
The results for the average duration of the computationally heavy linear regression step are shown in Figure 4. Minos lead to better linear regression times on all days the experiment was repeated, with different effect sizes. Over all days, Minos improved the average regression duration by $7.8\%$. This shows that Minos reaches its core goal, reducing the duration of the resource intensive computation.
@@
Figure 5. Successful requests per day. Minos improved the amount of successful requests all days except one with a maximum of $7.3\%$ on day one, but reduced the amount by $<1\%$ on day five. Note that the y-axis is limited to values between 4,000 and 5,000.
Extending the view from just looking at the resource intensive part of the function, Figure 5 shows how many functions are successfully executed by the ten VUs each day. On the first day, Minos manages to complete 7.3% more requests in the same 30 minutes. This drops in the next days, and at day five, the baseline completed slightly more requests. Over all experiments, Minos completed 2.3% more requests.
@@
Figure 6. Average Cost per million successful requests per day the experiment ran. On the first and last day, Minos saves more than $3\%$ cost, while it closely tracks the baseline on other days. Note that the y-axis is limited to values between 12 and 14.
We calculate the cost per invocation based on the cost model of Google Cloud Functions. For Minos, we also take the invocations that ultimately were terminated into account and calculate the average cost per actually completed request (successful request). As shown in Figure 6, on day one, Minos saves 3.3% of cost, while also having 3.4% lower latency. This small difference is due to the cost per invocation. On the other days, the cost per successful invocation of Minos is very close to the baseline. This shows that the platform itself is highly variable (as the experiment parameters never changed), and that the performance of Minos depends on the overall cloud platform performance. Over all experiments, Minos was able to reduce the cost per request by 0.9%.
Figure 7. Average cost per million successful requests over all experiments. While the difference is small, Minos has lower cost per successful request for 76% of the experiment duration. Note that the y-axis is limited to values between 10 and 25.
Looking at how the cost develops over the duration of the experiments (cf. Figure 7) shows that, at the beginning, Minos has a higher cost per successful request for the first 200s. This is expected, as Minos terminated many instances, increasing overall cost. Afterwards, the average cost per successful invocation stays quite close, with Minos staying slightly under the baseline for almost the whole duration. After 670 s (11 min 10 s), Minos was cheaper for more than 50 % of time, which increases to 76% at the end of the experiment. This shows that letting Minos run for a longer time increases its benefits.
4. Discussion & Future Work
@@
Future work could focus on researching other cost models that make wasting resources of the cloud provider cost prohibitive. One method could be billing users for the actual CPU performance delivered, not just the theoretical performance, to make less powerful instances financially applicable. This would however require platforms to make it explicit that users can not control the exact pricing and speed of their invocations. This is true today, it is just not represented in the cost model accordingly. Another method would be for platforms to switch to fairer virtualization technologies. For example, the popular FaaS platform AWS Lambda uses microVMs (Brooker et al., 2021), which in our preliminary experiments had a far fairer resource distribution with less overall variability.
@@
Our prototype for Minos pre-calculates the elysium threshold. It could also be updated live based on the benchmarking results of currently running instances. After a benchmark has finished, instances can report their results to a centralized component that collects live performance data. It then periodically re-calculates the elysium threshold and updates the functions comparable to our approach in our previous work (Schirmer et al., 2024). While this component is centralized, it is not a single point of failure as its failures would just lead to temporarily suboptimal performance. Depending on the scale of the application, it is not feasible to store all past benchmark results to accurately calculate the performance percentiles. Instead, some statistical properties can be calculated online, i.e., only having to store the most recent new value and a limited amount of information about past results. For example, the mean can be updated online by storing the previous mean and the amount of previous results. Other values, such as the standard deviation (Jain and Chlamtac, 1985) or percentiles (and, 1962) can only be estimated online and not calculated exactly.
@@
Note that Minos requires an asynchronous workload, i.e., users need to put their invocations into a queue instead of synchronously waiting for a response. When users require a synchronous response, the instance that receives the request needs to stay alive until the computation is finished to send the response back. If the instance fails the benchmark and should be crashed, it would need to send the call to another function and wait for the response, leading to double billing as the waiting instance still incurs cost (Baldini et al., 2017).
5. Related Work
Overall variability of FaaS functions has, e.g., been shown over longer (Eismann et al., 2022; Ustiugov et al., 2021) and shorter (Schirmer et al., 2023a) time frames, between processor architectures (Lambion et al., 2022), between platforms (Grambow et al., 2024), and for the overall platform behavior (Mahmoudi and Khazaei, 2021; Schirmer et al., 2023b; Ustiugov et al., 2021). Besides compute performance, instance placement can also be used to optimize for fewer carbon emissions (Chadha et al., 2023; Cordingly et al., 2023).
Ginzburg and Freedman present a method for using short-term intra-region FaaS instance variability on AWS Lambda (Ginzburg and Freedman, 2021). They demonstrate temporal, spatial (i.e., using other cloud regions), and instance performance variations in AWS Lambda in 2021. Their evaluation shows that exploiting instance performance variability can lead to performance increases. Compared to Minos, their approach focuses on demonstrating the existence of the effect for CPUs and relies on a centralized scheduler, only working for a limited amount of instances.
Another way of optimizing FaaS applications is speeding up cold starts, e.g., by pre-warming functions before they are called (Bermbach et al., 2020; Bardsley et al., 2018; Carl et al., 2024; Daw et al., 2020). This can be combined with Minos by benchmarking the pre-warmed instances before they are used. Other approaches focus on changing the underlying platform to speed up cold starts, e.g., by changing the virtualization and storage layer (Brooker et al., 2023), or by switching out the scheduling mechanism  (Cvetković et al., 2024).
Another approach is only using FaaS for the parts of the application where its most useful. Other parts can be executed to less expensive compute services (Zhao et al., 2023; Sadeghian et al., 2023).
6. Conclusion

### v2

# Trever Schirmer, Valentin Carl, Nils Höller, Tobias Pfandzelter, David Bermbach Technische Universität Berlin Scalable Software Systems Research Group

@@ Abstract
Index Terms:
Serverless, FaaS, workflows, resource contention
I Introduction
@@
Function-as-a-Service (FaaS) offerings have gained increasing popularity in the last years due to their ease of use [1, 2]. With FaaS, the cloud platform is responsible for handling most operational tasks, leaving users free to just write the code that should be executed for every incoming event, commonly referred to as functions [3, 1]. These functions can then be triggered by events in the cloud platform (such as a document upload) or by common transport protocols such as HTTP. In response to incoming events, the cloud platform assigns a function instance [4]. If there is no idle already running instance, the platform needs to start a new instance. This process, called a cold start, leads to increased latency for the request as a worker node needs to load the user code and start an isolated environment to run the user code in [5]. Often, platforms use microVMs [6] or containerization technologies for this [7]. Users are billed based on the execution time of the invocation, with millisecond accuracy [1]. This makes the model cost-efficient for use cases with irregular or infrequent usage patterns, since there is no resource overhead that users need to maintain.
Figure 1: FaaS platforms start new function instances on shared worker nodes. While users can not influence which worker node will be used, underutilized nodes offer two benefits: faster execution times and lower cost. Using the Minos system, newly started instances check if they are running on a node with low utilization. If not, they terminate themselves by crashing. This leads to a pool of better-performing instances that are re-used for subsequent invocations, leading to a compound performance increase.
As a cloud service hosted on shared infrastructure, functions suffer from performance variations between instances based on the usage patterns of the rest of the platform [2]. For example, as shown in Fig. 1, the worker node that the platform selects to start the instance can have different levels of overall utilization. The instance on the worker node with low utilization will benefit from having to share its resources with fewer neighbors. Taking the CPU as example this means that it will have to perform fewer context switches and that it is more likely data can remain cached. These variations also concern most other cloud services. However, FaaS users suffer twice when the variation leads to them having to use slower resources: not only is the execution time longer, but users are also billed more for these slower requests due to the pay-per-use model. In previous work [8], we have shown that some FaaS platforms have more than 10% faster performance during the night, when the overall platform is under less load. In this work, we focus on another kind of variation: when starting multiple instances in parallel, some of these instances are going to be faster than other ones. Selectively only running invocations on the fast instances leads to faster and cheaper function executions.
Figure 2: Overview of the process inside a function instance. The first step of the invocation, called prepare, is always executed. If the invocation is a cold start, a benchmark is executed in parallel. Minos then judges whether the result is better than the elysium threshold, in which case nothing happens (and the instance will be re-used in the future). If the instance does not pass, further execution is interrupted, the invocation is re-queued, and the instance is terminated.
To enable this, we present Minos, a system that runs a quick benchmark during every cold start. If the benchmark reveals a particularly slow instance, Minos re-queues the invocation and then terminates the function instance by crashing it. This leads to the platform starting a new instance that might have better performance, or scheduling the invocation on an already running instance. All instances that are already running must have passed the benchmark during their first request (as otherwise, they would have been shut down). This leads to a pool of known-good instances that outperform the average instance, which is especially useful for longer running applications that re-use these instances often. Our evaluation shows that this leads to a state where it is cheaper and faster for users to waste more resources of the platform.
@@
II A System to Exploit Platform Variability
@@
The following sections go into more detail on choosing how much should be terminated (Section II-A), how to decide on terminating in a distributed manner (Section II-B), and how the benchmark should work (Section II-C).
II-A How much to terminate?
@@
Figure 3: The total cost of a workflow comprises the cost per millisecond of execution ($c_{\textsf{exec}}$) of all terminated ($d_{\textsf{term}}$), successful ($d_{\textsf{pass}}$) benchmarks and the re-used instances ($d_{\textsf{reuse}}$), as well as the cost per invocation ($c_{\textsf{inv}}$) per terminated, passed, and re-used function.
As shown in Fig. 3, the total cost of a workflow is influenced by the cost per unit of execution and the cost per invocation. In general, the cost per unit of execution dominates the total cost: For the smallest Google Cloud Functions function with 128 MB Memory, the cost per invocation ($c_{\textsf{inv}}$) is roughly equivalent to 50ms of execution time ($c_{\textsf{exec}}$). For the biggest function with 32 GB Memory, it is less than 3 ms. Other platforms follow a similar pricing pattern. Thus, for all longer-running functions, the additional invocations made by Minos are quickly offset by using faster instances.
[…]
II-B Autonomous Termination with the Elysium Threshold
It would be optimal to have a centralized component that receives the benchmarking results from all instances to decide what instances are routed requests. While this would allow for complicated scheduling mechanisms, it is not feasible for the kind of highly scalable and elastic workloads that FaaS platforms enable users to run. Instead, a newly started function instance decides locally whether it is good enough based on just one value: what result in the benchmark is necessary to not terminate itself, i.e., so that the function will continue living and working on the next invocations. We call this value the elysium factor. This is stored as part of the function configuration, so that Minos does not require any outside communication during calls. Due to the short-lived nature of FaaS functions, re-running the benchmarks on an already accepted instance is not necessary, as instances are terminated after a short time.
@@ II-B Autonomous Termination with the Elysium Threshold
[…]
II-C What and how long to benchmark
The kind of benchmark that should be performed and the acceptable elysium threshold heavily depends on the use case. Previous research has shown that FaaS workloads are most often constrained by network or CPU resources [9]. In our preliminary experiments, CPU performance showed the highest variability and was the bottleneck in most applications.
The benchmark itself should ideally not interfere with the normal execution of the function, i.e., it should run while the resource being benchmarked is otherwise unused. The general approach is shown in Fig. 2. During a first step of the execution, when resources need to be set up and downloaded, the benchmark and subsequent judging is run in parallel. For example, in a data processing workflow, the raw data needs to be downloaded and loaded into memory. At the same time, Minos can run its benchmark.
@@ II-C What and how long to benchmark
III Evaluation
To demonstrate the feasibility of Minos and evaluate potential performance gains we implement a proof of concept prototype and use it to run a data processing workload. The prototype and dataset are available as open-source. [^2] https://github.com/nilshoeller/minos
[…]
III-A Environment Configuration
@@
The use case in this evaluation is using linear regression analysis to predict weather. To achieve this, the function first downloads a CSV file containing weather data for a specific location from previous days. Afterwards, the data is analyzed to predict tomorrow’s weather. While downloading the file, Minos can run its CPU benchmark as the download is network-limited. The following linear regression analysis is CPU intensive, benefitting from the instance selection performed in the previous step. To measure the capability of the CPU, we use matrix multiplication as benchmark in Minos [10].
@@
One experiment comprises ten virtual users (VU) that send a request, wait for it to complete, and then wait one more second before sending the next request over a total duration of 30 minutes. We repeat this experiment every day at the same time for one week. [^3] From 2025-02-03 to 2025-02-09, between 3 and 4 pm UTC As a baseline, we run the same workload at the same time using another function that is exactly the same, except that all components of Minos are disabled. All experiments were executed in the europe-west-3 region of Google Cloud Platform with functions configured to have 256 MB RAM (leading to 0.167 vCPU).
@@
For both the baseline and experiment conditions, we measure the time of the function execution, how long the downloading step took (download duration), and how long the linear regression took (analysis duration). Additionally, for the experiment condition, we measure the duration of the benchmark execution, whether it was successful, and how often the invocation has been re-tried (retry count).
III-B Results
@@
[…]
@@
Extending the view from just looking at the resource intensive part of the function, Fig. 5 shows how many functions are successfully executed by the ten VUs each day. On the first day, Minos manages to complete 7.3% more requests in the same 30 minutes. This drops in the next days, and at day five, the baseline completed slightly more requests. Over all experiments, Minos completed 2.3% more requests.
@@
We calculate the cost per invocation based on the cost model of Google Cloud Functions. For Minos, we also take the invocations that ultimately were terminated into account and calculate the average cost per actually completed request (successful request). As shown in Fig. 6, on day one, Minos saves 3.3% of cost, while also having 3.4% lower latency. This small difference is due to the cost per invocation. On the other days, the cost per successful invocation of Minos is very close to the baseline. This shows that the platform itself is highly variable (as the experiment parameters never changed), and that the performance of Minos depends on the overall cloud platform performance. Over all experiments, Minos was able to reduce the cost per request by 0.9%.
Figure 7: Average total cost per million successful requests over all experiments. While the difference is small, Minos has lower cost per successful request for 76% of the experiment duration. Note that the y-axis is limited to values between 10 and 25.
Looking at how the cost develops over the duration of the experiments (cf. Fig. 7) shows that, at the beginning, Minos has a higher cost per successful request for the first 200s. This is expected, as Minos terminated many instances, increasing overall cost. Afterwards, the average cost per successful invocation stays quite close, with Minos staying slightly under the baseline for almost the whole duration. After 670 s (11 min 10 s), Minos was cheaper for more than 50 % of time, which increases to 76% at the end of the experiment. This shows that letting Minos run for a longer time increases its benefits.
IV Discussion & Future Work
@@
Future work could focus on researching other cost models that make wasting resources of the cloud provider cost prohibitive. One method could be billing users for the actual CPU performance delivered, not just the theoretical performance, to make less powerful instances financially applicable. This would however require platforms to make it explicit that users can not control the exact pricing and speed of their invocations. This is true today, it is just not represented in the cost model accordingly. Another method would be for platforms to switch to fairer virtualization technologies. For example, the popular FaaS platform AWS Lambda uses microVMs [6], which in our preliminary experiments had a far fairer resource distribution with less overall variability.
@@
Our prototype for Minos pre-calculates the elysium threshold. It could also be updated live based on the benchmarking results of currently running instances. After a benchmark has finished, instances can report their results to a centralized component that collects live performance data. It then periodically re-calculates the elysium threshold and updates the functions comparable to our approach in our previous work [11]. While this component is centralized, it is not a single point of failure as its failures would just lead to temporarily suboptimal performance. Depending on the scale of the application, it is not feasible to store all past benchmark results to accurately calculate the performance percentiles. Instead, some statistical properties can be calculated online, i.e., only having to store the most recent new value and a limited amount of information about past results. For example, the mean can be updated online by storing the previous mean and the amount of previous results. Other values, such as the standard deviation [12] or percentiles [13] can only be estimated online and not calculated exactly.
@@
Note that Minos requires an asynchronous workload, i.e., users need to put their invocations into a queue instead of synchronously waiting for a response. When users require a synchronous response, the instance that receives the request needs to stay alive until the computation is finished to send the response back. If the instance fails the benchmark and should be crashed, it would need to send the call to another function and wait for the response, leading to double billing as the waiting instance still incurs cost [14]. Additionally, the functions need to have an initial phase where the resource that should be benchmarked is not the bottleneck and need to take long-enough that running a benchmark at all is feasible. While this is true for common workloads in, e.g., machine learning inference, it is not universally applicable.
V Related Work
Overall variability of FaaS functions has, e.g., been shown over longer [15, 16] and shorter [17] time frames, between processor architectures [18], between platforms [19], and for the overall platform behavior [20, 8, 16]. Besides compute performance, instance placement can also be used to optimize for fewer carbon emissions [21, 22].
Ginzburg and Freedman present a method for using short-term intra-region FaaS instance variability on AWS Lambda [23]. They demonstrate temporal, spatial (i.e., using other cloud regions), and instance performance variations in AWS Lambda in 2021. Their evaluation shows that exploiting instance performance variability can lead to performance increases. Compared to Minos, their approach focuses on demonstrating the existence of the effect for CPUs and relies on a centralized scheduler, only working for a limited amount of instances.
Another way of optimizing FaaS applications is speeding up cold starts, e.g., by pre-warming functions before they are called [24, 25, 26, 27]. This can be combined with Minos by benchmarking the pre-warmed instances before they are used. Other approaches focus on changing the underlying platform to speed up cold starts, e.g., by changing the virtualization and storage layer [4], or by switching out the scheduling mechanism  [28].
Another approach is only using FaaS for the parts of the application where its most useful. Other parts can be executed to less expensive compute services [29, 30].
VI Conclusion